OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (2, -5)with a slope of -3. y = -3x - 13 y = -3x + 1 y = -3x + 13 y = -3x + 11

OpenStudy (anonymous):

(y-(-5))/(x-2) = -3 (y+5)/(x-2) = -3 y+5 = -3(x-2) y+5 = -3x+6 y=-3x+1 the second one is the answer

OpenStudy (anonymous):

Thanks again ^_^